# Hybrid Quantum-inspired Resnet and Densenet for Pattern Recognition

### v6

@@ 5.2.1 Test under symmetrical noise attacking
mixed 0.8879 0.6839 0.9075
mixed 0.8326 0.8242 0.9130
gaussian 0.8634 0.8792 0.9109
gaussian 0.8701 0.8893 0.9267
uniform 0.8943 0.8742 0.9214
uniform 0.9045 0.9001 0.9148
Fig.5 and Table 10 demonstrate the accuracy of HQNet, QRCNN and QDCNN on FASHIONMNIST under symmetrical attacks in the two forms. From the outcomes, QRCNN does not outperform HQNet, but QDCNN directy shows more advantages with the densely-connected architecture over QRFNN and HQNet. For instance, under mixed noise attacking in form ②, while the accuracy of QRCNN and HQNet is 83.26% and 82.42%, that of QDCNN is 91.30%, which is 8%-9% higher than that of HQNet and QRCNN. Comparison results between pure classical models and our hybrid models are in the part 8 of the supplementary material. Sometimes our hybrid models show slight more prominent performance than classical models, while the results of classical models are marginally better than that of our hybrid models at times. On the whole, our hybrid models indicate roughly the same level of capability as classical models in resistance to symmetrical noise attacks.
5.2.2 Test under unsymmetrical noise attacking
Figure 6: Unsymmetrical noise attack in form ① in CIFAR100. We utilize two pure classical CNNs for comparison, of which the activation functions are separately rot-relu and leaky-relu functions. (a) Parameter attack with mixed unsymmetrical noise. (b) Parameter attack with gaussian unsymmetrical noise. In terms of the average test accuracy of the two pure classical models, the results remain low values in (a) and (b) since the loss reach nan.
Figure 7: Accuracy under unsymmetrical noise attack in form ② on FASHIONMNIST. We utilize two pure classical CNNs for comparison, of which the activation functions are separately rot-relu and leaky-relu function. (a) Parameter attack with mixed unsymmetrical noise. (b) Parameter attack with gaussian unsymmetrical noise. In terms of the average test accuracy of the two pure classical models, the results remain low values in (a) and (b) since the loss reach nan.
Figure 8: Accuracy under unsymmetrical mixed noise attacks on FASHIONMNIST. We utilize a state-of-the-art HQNet for comparison. (a) Parameter attack in form ①. (b) Parameter attack in form ②.
Table 11: Accuracy on FASHIONMNIST under unsymmetrical noise attacking
@@ 5.2.2 Test under unsymmetrical noise attacking
Table 12: Accuracy on FASHIONMNIST under unsymmetrical noise attacking
@@
mixed 0.8278 0.8401 0.8999
mixed 0.6833 0.7380 0.9187
gaussian 0.7993 0.7612 0.9056
gaussian 0.8596 0.8478 0.9125
uniform 0.7934 0.8150 0.9073
uniform 0.8092 0.8445 0.9177
More importantly, our hybrid models show much greater superiority over classical models under unsymmetrical noise attacks in the two forms (see supplementary materials part 8 for more results). Fig.6 shows the accuracy under gaussian and mixed asymmetric attack in form ① on CIFAR100. As for Fig.(6a), while the accuracy of two CNNs is only 25% and 23%, that of QDCNN and QRCNN is 78.67% and 73.65% respectively, which shows the advantages of the quantum-inspired layers of our hybrid architectures. And there are also a little more superiority of QDCNN over QRCNN. The accuracy results of the four models in Fig.(6b) is similar with that in Fig.(6a). Fig.7 demonstrates the accuracy under gaussian and mixed asymmetric attacks in form ② on FASHIONMNIST. In the two figures, the accuracy of the pure classical models maintains at a low value around 25%, which shows that the pure classical models fail to learn the features. And we find the loss curves of the pure classical models oscillate badly and the loss values are very large and up to nan, which implies the reason of the bad performance of the pure classical models is gradient explosion (see part 9 of the supplementary material). However, even under various unsymmetrical noise attacks with the two forms, our hybrid models still show great performance on the tasks with the accuracy being 90% roughly due to the quantum-inspired layers in our hybrid models to avoid gradient explosion. And Table 11 shows the average accuracy values of the curves in Fig.6 and Fig.7.
@@ 5.2.3 Ablation study
To test the advantages of the adaptive residual and dense connection of our hybrid models, we use Traditional Quantum-inspired Feedforward Neural Network (TQFNN) and Traditional Quantum-inspired Convolutional Neural Network (TQCNN) for comparison. The framework of TQFNN is exactly the same as that of QRFNN and QDFNN but without residual or dense connection. Similarly, the architecture of TQCNN is also identical to that of QRCNN and QDCNN but without residual or dense connection in the convolutional and fully connected layers as well. Hence, through theoretical analysis in the part 4 of the supplementary material, the TQFNN, QRFNN and QDFNN models are close in the computational and parameter complexity. Moreover, experimental results demonstrate that the accuracy of our hybrid models is approximately 2%-3% higher than that of TQFNN and TQCNN. For example, as for Fig.(9a), the accuracy of TQFNN on clean iris data is 87.93%, while that of QRFNN and QDFNN is 89.57% and 92.55%. It indicates the superiority of the adaptive residual and dense connections in our hybrid models over the quantum-inspired neural networks without residual or dense connection. Furthermore, QDFNN and QDCNN outperform QRFNN and QRCNN a little separately, which shows the more advantages of the dense connection in QDFNN and QDCNN over the residual connection in QRFNN and QRCNN.
Figure 9: Test accuracy of the traditional quantum-inspired neural network and our hybrid models with noiseless datasets. (a) Accuracy results of noiseless iris data of TQFNN, QRFNN and QDFNN. (b) Accuracy results of noiseless FASHIONMNIST data of TQCNN, QRCNN and QDCNN. The average accuracy of TQCNN is 90.19%, while that of QRCNN and QDCNN is 92.52% and 93.55%. It indicates the slightly more advantage of the dense connection in QDCNN over the residual connection in QRCNN. Additionally, our two hybrid models outperform TQCNN, which shows the function of the residual and dense connections.
@@ 6 Conclusion and discussion
To summarize, we have firstly proposed the hybrid quantum-inspired neural networks combining symmetrical circuit model with adaptive residual or dense connections. We explain their frameworks and assess their generalization power and robustness. HQNet and Classical MLP, CNNs with concrete structures and quantum-inspired neural networks without residual or dense connection are used for comparison. The charm of our hybrid models lies in these facts that:
@@
    They show much more outstanding robustness than the MLPs and CNNs with the parameters attacked by unsymmetrical noises;
@@
    The densely-connected frameworks are more advantageous than HQNet on classification tasks under the attacks of symmetrical and asymmetrical noises;
@@
    They raise potential to systematically prevent the gradient explosion problem.
While the quantum-inspired layers enhance the robustness and prevent gradient explosion, the adaptive residual and dense connections facilitate feature learning and generalization power of our hybrid models. Therefore, the novelty of our models lies in the combination of adaptive residual or dense connection with the quantum-inspired layers with symmetrical circuits, which improves the comprehensive performance in the neural network field. Hence, our hybrid models are able to substitute pure classical or other quantum-inspired neural networks in certain tasks. However, according to the part 5 of the supplementary material, the computational complexity of QRFNN, $C_{QRFNN}\approx B\cdot I\cdot L\cdot N^{4}\propto N^{4}$, and the computational complexity of QDFNN, $C_{QDFNN}\approx B\cdot I\cdot L\cdot N^{4}\propto N^{4}$, where $B$ and $I$ are batchsize and number of iterations in an epoch separately. Since $N=2^{n}$, $C_{QRFNN}\propto 2^{4n}$, and $C_{QDFNN}\propto 2^{4n}$, where n is the number of qubit. It indicates the trouble to expand the dimension of the quantum state in the quantum-inspired layer, which is also the width of our hybrid neural networks [19, 51]. If our hybrid models are used for large-scale datasets, the computational complexity of them could will increase exponentially, leading to a large amount of time consumption. Therefore, our models are more suitable for problems on data of modest-scale dimension of the feature space, such as iris data in scikit-learn library [31]. Nonetheless, when it comes to the design of deep neural networks, our work may herald a future where the hybrid architectures redefine the boundaries of deep learning to unprecedented heights. What’s more, owing to the outstanding generalization capability and robustness of our hybrid models, it makes sense to leverage them on complicated noisy engineering environments [47]. The links of the supplementary material and program codes are below the acknowledgements.
CRediT authorship contribution statement
Andi Chen: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. Hua-Lei Yin: Writing – review & editing. Zeng-Bing Chen: Writing – review & editing. Shengjun Wu: Writing – review & editing.

### v7

@@ 5.2.1 Test under symmetrical noise attacks
mixed 0.8879$\pm$0.0102 0.6839$\pm$0.0132 0.9075$\pm$0.0096
mixed 0.8326$\pm$0.0116 0.8242$\pm$0.0117 0.9130$\pm$0.0094
gaussian 0.8634$\pm$0.0109 0.8792$\pm$0.0103 0.9109$\pm$0.0096
gaussian 0.8701$\pm$0.0108 0.8893$\pm$0.0102 0.9267$\pm$0.0091
uniform 0.8943$\pm$0.0105 0.8742$\pm$0.0110 0.9214$\pm$0.0091
uniform 0.9045$\pm$0.0100 0.9001$\pm$0.0099 0.9148$\pm$0.0088
Fig.5 and Table 12 demonstrate the accuracy of HQNet, QRCNN and QDCNN on FASHIONMNIST under symmetrical attacks in the two forms. From the outcomes, QRCNN does not outperform HQNet, but QDCNN directy shows more advantages with the densely-connected architecture over QRCNN and HQNet. For instance, under mixed noise attacking in form ②, while the accuracy of QRCNN and HQNet is 83.26%$\pm$1.16and 82.42%$\pm$1.17%, that of QDCNN is 91.30%$\pm$0.94, which is 8%-9% higher than that of HQNet and QRCNN. Comparison results between pure classical models and our hybrid models are in the section 8 of the supplementary material. Sometimes our hybrid models show slight more prominent performance than classical models, while the results of classical models are marginally better than that of our hybrid models at times. On the whole, our hybrid models indicate roughly the same level of capability as classical models in resistance to symmetrical noise attacks.
5.2.2 Test under asymmetrical noise attacks
Figure 6: Asymmetrical noise attack in form ① in CIFAR100. We utilize two pure classical CNNs for comparison, of which the activation functions are separately rot-relu and leaky-relu functions. (a) Parameter attack with mixed asymmetrical noise. (b) Parameter attack with gaussian asymmetrical noise. In terms of the average test accuracy of the two pure classical models, the results remain low values in (a) and (b) since the loss reach nan.
Figure 7: Accuracy under asymmetrical noise attack in form ② on FASHIONMNIST. We utilize two pure classical CNNs for comparison, of which the activation functions are separately rot-relu and leaky-relu function. (a) Parameter attack with mixed asymmetrical noise. (b) Parameter attack with gaussian asymmetrical noise. In terms of the average test accuracy of the two pure classical models, the results remain low values in (a) and (b) since the loss reach nan.
Figure 8: Log values of gradient of an attacked parameter. The noise utilized is asymmetrical gaussian one. As the graph shows, since gradient explosion occurs in two CNNs, their gradient log values increase exponentially, while those of QRCNN and QDCNN are close to 0.
Figure 9: Accuracy under asymmetrical mixed noise attacks on FASHIONMNIST. We utilize a state-of-the-art HQNet for comparison. (a) Parameter attack in form ①. (b) Parameter attack in form ②.
Table 13: Accuracy on FASHIONMNIST under asymmetrical noise attacking
@@ 5.2.2 Test under asymmetrical noise attacks
Table 14: Accuracy on FASHIONMNIST under three asymmetrical noise attacks
@@
mixed 0.8278$\pm$0.0107 0.8401$\pm$0.0102 0.8999$\pm$0.0092
mixed 0.6833$\pm$0.0122 0.7380$\pm$0.0117 0.9187$\pm$0.0087
gaussian 0.7993$\pm$0.0123 0.7612$\pm$0.0128 0.9056$\pm$0.0091
gaussian 0.8596$\pm$0.0099 0.8478$\pm$0.0100 0.9125$\pm$0.0088
uniform 0.7934$\pm$0.0122 0.8150$\pm$0.0111 0.9073$\pm$0.0095
uniform 0.8092$\pm$0.0110 0.8445$\pm$0.0116 0.9177$\pm$0.0087
More importantly, our hybrid models show much greater superiority over classical models under asymmetrical noise attacks in the two forms (see supplementary materials section 8 for more results). Fig.6 shows the accuracy under gaussian and mixed asymmetrical attack in form ① on CIFAR100. As for Fig.(6a), while the accuracy of two CNNs is only 25%$\pm$1.45and 23%$\pm$1.43separately, that of QDCNN and QRCNN is 78.67%$\pm$1.23and 73.65%$\pm$1.29respectively, which shows the advantages of the quantum-inspired layers of our hybrid architectures. And there are also a little more superiority of QDCNN over QRCNN. The accuracy results of the four models in Fig.(6b) is similar with that in Fig.(6a). Fig.7 demonstrates the accuracy under gaussian and mixed asymmetrical attacks in form ② on FASHIONMNIST. In the two figures, the accuracy of the pure classical models maintains at a low value around 25%, which shows that the pure classical models fail to learn the features. And we find the loss curves of the pure classical models oscillate badly and the loss values are very large and up to nan, which implies the reason of the bad performance of the pure classical models is gradient explosion (see section 9 of the supplementary material). However, even under various asymmetrical noise attacks with the two forms, our hybrid models still show great performance on the tasks with the accuracy being 90% roughly due to the quantum-inspired layers in our hybrid models to avoid gradient explosion. And Table 13 shows the average accuracy values of the curves in Fig.6 and Fig.7.
To further validate whether the reason of the terrible performance of classical models is from gradient explosion, we print the logarithm values of the gradient of the loss function to an attacked parameter of our hybrid models and classical CNNs. We find that the gradient values of CNNs grow rapidly and reach nan with epoch increasing. Hence, we have to print logarithm values of the gradient of an attacked parameter of CNNs. As is shown in Fig.8, while the curves of QRCNN and QDCNN are flat, of which the values are close to 0, those of two CNNs rise rapidly, which is consistent with the characteristics of gradient explosion.
@@
Figure 10: Accuracy under asymmetrical noise attacks on CIFAR10 in form ①. We utilize ResQNet1 and ResQNet2 for comparison. (a) uniform noise attack (b) mixed noise attack.
Figure 11: Accuracy under asymmetrical noise attacks on CIFAR10 in form ②. We utilize ResQNet1 and ResQNet2 for comparison. (a) gaussian noise attack (b) mixed noise attack.
Table 15: Accuracy on CIFAR10 under three asymmetrical noises attack in form ① and ②
attack form noise type ResQNet1 ResQNet2 QRCNN QDCNN
uniform 76.21%$\pm$2.0274.55%$\pm$2.0777.95%$\pm$1.9777.92%$\pm$1.97%
mixed 78.80%$\pm$1.9477.84%$\pm$1.9779.81%$\pm$1.9178.80%$\pm$1.95%
mixed 76.47%$\pm$2.0278.20%$\pm$1.9678.07%$\pm$1.9780.31%$\pm$1.89%
gaussian 77.41%$\pm$1.9975.79%$\pm$2.0377.20%$\pm$1.9979.20%$\pm$1.93%
Besides that, Fig.10 shows the accuracy curves of ResQNet1, ResQNet2 and our hybrid models to three asymmetrical noise attacks in form ① on CIFAR10. And Fig.11 shows that of ResQNet1, ResQNet2 and our hybrid models to three asymmetrical noise attacks in form ② on CIFAR10. All the curves in both the figures firstly rise and then fluctuate around certain values. Also, Table 15 showcases the exact average values of accuracy of the four models with standard deviation in Fig.10 and Fig.11, from which we can see QRCNN and QDCNN slightly outperform ResQNet1 and ResQNet2. For instance, shown on Table 15, under gaussian noise attacks in form ②, while the accuracy of ResQNet1 and ResQNet2 on CIFAR10 is separately 77.41% and 75.79%, that of QRCNN and QDCNN is respectively 77.20% and 79.20%. It also tells the advantage of the residual and dense mappings in our hybrid models. It is also valuable to explore the combination of circuits and other residual and dense connections to further facilitate the accuracy of hybrid neural networks.
@@ 5.2.3 Ablation study
To test the advantages of the adaptive residual and dense connection of our hybrid models, we use Traditional Quantum-inspired Feedforward Neural Network (TQFNN) and Traditional Quantum-inspired Convolutional Neural Network (TQCNN) for comparison. The framework of TQFNN is exactly the same as that of QRFNN and QDFNN but without residual or dense connection. Similarly, the architecture of TQCNN is also identical to that of QRCNN and QDCNN but without residual or dense connection in the convolutional and fully-connected layers as well. Hence, through theoretical analysis in the section 4 of the supplementary material, the TQFNN, QRFNN and QDFNN models are close in the computational and parameter complexity. Moreover, experimental results demonstrate that the accuracy of our hybrid models is approximately 2%-3% higher than that of TQFNN and TQCNN. For example, as for Fig.(12a), the accuracy of TQFNN on clean iris data is 87.93%$\pm$1.04%, while that of QRFNN and QDFNN is 89.57%$\pm$0.98and 92.55%$\pm$0.93%. It showcases the superiority of the adaptive residual and dense connections in our hybrid models over the quantum-inspired neural networks without residual or dense connection. Furthermore, QDFNN and QDCNN outperform QRFNN and QRCNN a little separately, which shows the more advantages of the dense connection in QDFNN and QDCNN over the residual connection in QRFNN and QRCNN.
Figure 12: Test accuracy of the traditional quantum-inspired neural network and our hybrid models with noiseless datasets. (a) Accuracy results of noiseless iris data of TQFNN, QRFNN and QDFNN. (b) Accuracy results of noiseless FASHIONMNIST data of TQCNN, QRCNN and QDCNN. The average accuracy of TQCNN is 90.19%$\pm$0.95%, while that of QRCNN and QDCNN is 92.52%$\pm$0.91and 93.55%$\pm$0.86%. It indicates the slightly more advantage of the dense connection in QDCNN over the residual connection in QRCNN. Additionally, our two hybrid models outperform TQCNN, which shows the function of the residual and dense connections.
@@ 6 Conclusion and discussion
To summarize, we have firstly proposed the hybrid quantum-inspired neural networks amalgamating symmetrical circuits with adaptive residual or dense connections. We explain their frameworks and assess their generalization power and robustness. ResQNet1, ResQNet2, HQNet and classical MLP, CNNs with concrete structures and traditional quantum-inspired neural networks without residual or dense connection are for comparison. The charm of our hybrid models lies in these facts that:
@@
    they show much more remarkable robustness than the MLPs and CNNs with the parameters attacked by asymmetrical noises;
@@
    the densely-connected frameworks are a bit more advantageous than HQNet on classification tasks under the attacks of symmetrical and asymmetrical noises;
@@
    their frameworks are a little bit more superior than ResQNet1 and ResQNet2 on classification tasks with asymmetrically noisy datasets or to asymmetrical noise attacks;
  • 7.
    They raise great potential to systematically prevent the gradient explosion problem.
While the quantum-inspired layers enhance the robustness and prevent gradient explosion, the adaptive residual and dense connections facilitate feature learning and generalization power of our hybrid models. Therefore, the biggest novelty of our models lies in the combination of the proposed adaptive residual or dense connection with proposed symmetrical circuits, which improves the comprehensive performance in the neural network field. Hence, our hybrid models are able to substitute pure classical or other quantum-inspired neural networks in certain tasks. However, according to the section 5 of the supplementary material, the computational complexity of QRFNN, $C_{QRFNN}\approx B\cdot I\cdot L\cdot N^{4}\propto N^{4}$, and the computational complexity of QDFNN, $C_{QDFNN}\approx B\cdot I\cdot L\cdot N^{4}\propto N^{4}$, where $B$ and $I$ are batchsize and number of iterations in an epoch separately. Since $N=2^{n}$, $C_{QRFNN}\propto 2^{4n}$, and $C_{QDFNN}\propto 2^{4n}$, where n is the number of qubit. It indicates the trouble to expand the dimension of the quantum state in the quantum-inspired layer, which is also the width of our hybrid neural networks [19, 52]. If our hybrid models are utilized for large-scale datasets, the computational complexity of them could increase exponentially, leading to a large amount of time consumption. Therefore, our models are more suitable for problems on data of modest-scale dimension of the feature space, such as iris data in scikit-learn library [31]. Nonetheless, when it comes to the design of deep neural networks, our work may herald a future where the hybrid architectures redefine the boundaries of deep learning to unprecedented heights. What’s more, owing to the outstanding generalization capability and robustness of our hybrid models, it makes sense to leverage them on complicated noisy engineering environments involving the coupled effects of multiple factors, such as unstructured data and noise interference [48]. To exemplify, it does make sense to evaluate the robustness of our hybrid models when there exists a substantial distributional discrepancy between training data and practical engineering scenarios, which would significantly facilitate the applicability of our models in real-world contexts. The links of the supplementary material and program codes are below the acknowledgements.